OpenStudy (anonymous):

What is the solution of the equation 6x-3=-51? A. -9 B.-8 C. 8 D. 9

OpenStudy (anonymous):

which do you think it is mate?

OpenStudy (anonymous):

lets start with A: 6(-9)-3=51 would be FALSE because -9*6=-54 and -54-3= -57 so that's not it now B: 6(-8)-3=51 would be FALSE because -8*6=-48 and -48-3=-51 so that cant be right C: 6(8)-3=51 would be FALSE because 6*8= 48 and 48-3 would be 45 so that's definiftly not it D: 6(9)-3=51 would be TRUE because 6*9=54 and 54-3 would be 51 so that's right! so D would be your answer :)